The sum of two numbers is 57. One of the numbers is 21. Which of the following equations could be used to find the other number?
An answer is required
n = 57 + 21
57 = 2 + 21
21 = n – 57
Not Here
57 = n + 21

Answer by stanbon(75887)

The sum of two numbers is 57. One of the numbers is 21. Which of the following equations could be used to find the other number?
An answer is required
-------
Since n+21 = 57
n = 57-21
or
n-57 = -21
-----------------------
Cheers,
Stan H.
===================
n = 57 + 21
57 = 2 + 21
21 = n – 57
Not Here
57 = n + 21